A student has an average of 72 on four exams. If the final exam is to count double, what score does she need to have an averge of 80?
Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

(4*72+2x)/6=80
x=96
at least a score of 96
another way
72 misses 80 by 8
4*8=32
32/2=16
16+80=96


Answer by solver91311(24713)



If his current average is 72 for 4 exams, then his total points have to be 4 times 72.

If his final average for 4 exams plus a final that counts for 2 exams must be 80, then his final point total must be 6 times 80.

Subtract his current point total from his needed final total giving you the points he needs to achieve. Since the final counts as 2, this number will be 2 times the score he needs on the final.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism